OpenStudy (anonymous):

7/8w = - 2/3 What are the steps to solve.

OpenStudy (anonymous):

does it read \[\frac{ 7 }{ 8 }w\] or \[\frac{ 7 }{ 8w }\]?

OpenStudy (anonymous):

8*(7/8)w = (-2/3)*8 7w = (-2/3)*8 (7w)/7=(-2/3)*(8/7) w= -16/21